Find two numbers whose sum is 55 and whose product is 684.
This is my homework. But I would really appreciate it if you could write the formula so that I can also understand how you did it and can further use it in my future encounter with these problem.

Answer by Cromlix(4381)

x + y = 55....1
xy = 684......2
Alter (2)
x = 684/y......2
Substitute x = 684/y into (1)
x + y = 55
684/y + y = 55
Common Denominator = y
684 + y^2 = 55y
Rearrange
y^2 - 55y + 684 = 0
(y - 36)(y - 19) = 0
So y can equal 36 or 19
and x can equal 19 or 36.
Hope this helps.
:-)